OpenStudy (anonymous):

The second angle of a triangle is 25 more than twice the measure of the first angle. The third angle of a triangle is 15 degrees more than half the measure of the first angle. Find the measure of each angle. Can someone please show me a drawing of what this would look like?

OpenStudy (anonymous):

We're going to represent the first angle with "a". Since the second and third angles are defined by a, we can also represent these angles using this variable. the second angle = 2a+25 the third angle = a/2+15 since the total number of degrees in a triangle is 180 a+2a+25+a/2+15=180 combine like terms 3a+a/2+40=180 3a+a/2=140 (we subtracted 40 from both sides) 6a+a=280 (we multiplied both sides by 2 to remove the /2 from the second "a") 7a=280 a=40 (we divided both sides by 7 to remove the coefficient) now that we have found a, we can define the other angles with this value angle 2 = 2(40)+25 = 80+25 = 105 angle 3 = 40/2+15 = 20+15 = 35 we can check our work by adding all the angles together 40+105+35 = 40+140 = 180 this triangle would look something like this 

OpenStudy (anonymous):

okay thanks